The smaller of two consecutive even integers is 2 more than twice the larger. Find the numbers.
Answer by checkley77(12569)

x & (x+2) are the 2 even integers.
x=2(x+2)+2
x=2x+4+2
x-2x=6
-x=6
x=-6 answer for the smaller integer.
-6+2=-4 answer for the larger integer.
Proof:
-6=2*-4+2
-6=-8+2
-6=-6